# LLM-Guided Scenario-based GUI Testing

### v2

@@ 2.1. Testing Scenario
In real-world app usage, end users interact with mobile apps through a series of GUI operations, such as tapping, swiping, or text input, to accomplish a specific goal. These operation sequences form logical patterns that can be abstracted into corresponding testing scenarios. For example, the “user login” scenario may include operations like entering credentials, tapping the login button, and verifying access to the home page. Understanding such scenarios is critical for testing app functionality from a user-centric perspective, as it enables testers to validate whether the app behaves as intended under realistic usage contexts. In the practice of manual GUI testing, developers and testers typically organize test cases around distinct testing scenarios. This organization allows them to design test cases more systematically, ensuring that all major functionalities and user flows are adequately exercised. Testing scenarios thus serve as a bridge between app functionality and user behavior, facilitating a more intuitive and comprehensive verification of the system’s behavior.
@@ 3. Methodology
Specifically, the Observer perceives and interprets the app GUI by detecting widgets and extracting their semantic attributes, enabling the system to understand the current interface context. Building upon this perception, the Decider performs reasoning and planning, leveraging LLM capabilities to align GUI interactions with the target testing scenario and business logic. The Executor then executes the planned operations on the app, ensuring repeatability and traceability while decoupling action execution from decision-making to reduce the risk of cascading failures. Next, the Supervisor verifies whether the executed operations achieve the intended scenario objectives and provides feedback for correction when discrepancies occur. Finally, the Recorder records all operations, GUI states, and inspection outcomes into the shared context memory, preserving contextual information to support consistent and informed decision-making in subsequent iterations.
@@ 3.1. Context Memory
Due to the stateless nature of LLMs, it is essential to design an explicit mechanism for managing contextual information throughout the iterative testing process. To ensure that ScenGen can maintain continuity across iterations and generate coherent scenario-based GUI tests, we introduce a structured context memory inspired by human memory theory (Baddeley, 1992). The context memory serves as a central repository that persistently stores and organizes diverse types of information generated during testing, supporting reasoning, decision-making, and feedback across all agents. Drawing from the cognitive model of human memory, the context memory is divided into three components: long-term memory, working memory, and short-term memory, each serving a distinct purpose in the testing workflow.
@@ 3.2. Observer
The Observer is responsible for perceiving the app GUI and constructing a structured representation of its current state. This process combines traditional computer vision (CV) techniques with multi-modal LLM capabilities to achieve both precise widget localization and semantic understanding. Such a hybrid design leverages the strengths of CV in pixel-level perception and the reasoning capability of LLMs in contextual interpretation, thereby enabling accurate and semantically rich GUI state analysis.
@@ 3.3. Decider
Scenario-based GUI test generation in ScenGen is modeled as an iterative decision-making process, starting from the app’s initial state and progressing through a sequence of actions to cover a predefined target testing scenario. The process begins with a scenario prompt, which provides contextual guidance to the multi-agent system. Each decision iteration evaluates the current app state, reasons about the most appropriate next operation, and executes it until the scenario goal is fully achieved. Every action decision is derived as an inference mapping generated by an LLM, formulated as follows:
@@
Compared to traditional LLMs, multi-modal LLMs can directly process GUI screenshots, enabling visual-semantic reasoning and action planning. Thus, ScenGen employs raw GUI images as representations of app states and utilizes multi-modal LLMs to make visual decisions directly. However, this design introduces a cognition–execution gap: while LLMs excel at reasoning about what should be done, they lack the ability to locate precise pixel-level widget positions required for action execution (Liu et al., 2024b). To address this, ScenGen integrates multi-modal reasoning with traditional vision-based widget localization. The decision process is therefore divided into two stages, logical decision-making and widget localization, bridging high-level reasoning and low-level action feasibility.
@@ 3.3.2. Target Widget Location
After logical decision-making, the action intent and target description are available, but precise pixel-level widget coordinates are still unknown. To enable actionable execution, the Decider performs widget localization, refining the decision into an executable operation:
@@ 3.5. Supervisor
After each action decision is executed, the app state must be verified before the next decision to ensure that the execution outcome aligns with expectations. To achieve comprehensive and reliable validation, the Supervisor performs a three-stage verification process: real-time loading verification, state transition verification, and test completion verification. Together, these stages ensure that the app state is correctly recognized, that the system responds appropriately to both expected and unexpected outcomes, and that the testing process proceeds in accordance with the intended scenario.
@@
Real-time Loading Verification. The execution of a test operation may temporarily place the app into an unstable state, such as a loading or transition phase triggered by a network request. In such cases, premature decision-making may lead to misinterpretation of the app state or the early triggering of self-correction. While a fixed waiting time could mitigate this issue, it is difficult to generalize and often harms testing efficiency. Instead, ScenGen leverages the visual understanding capability of a multi-modal LLM to dynamically determine whether the app has completed loading. By analyzing the current GUI image, the Supervisor detects loading indicators or transitional states and adjusts the waiting duration adaptively. This adaptive waiting strategy improves both the reliability and efficiency of the testing process. An example of this verification is shown in Fig. 4 (a).
@@ 4.1. Experiment Settings
  • RQ1: How effective is the logical decision-making of ScenGen?
  • RQ2: How effective is the widget localization of ScenGen?
  • RQ3: How effective is the scenario-based test generation of ScenGen?
  • RQ4: How efficient is the scenario-based test generation of ScenGen?
  • RQ5: How effective is the bug detection capability of ScenGen?
These research questions jointly provide a comprehensive assessment of ScenGen from multiple perspectives. Specifically, RQ1 evaluates the accuracy of logical decision-making by the Decider, assessing the LLM’s ability to produce semantically valid and contextually relevant actions. RQ2 measures the precision of widget localization conducted by the Observer, focusing on accurate visual grounding within GUI states. RQ3 investigates the overall quality of the generated tests, verifying the ability of ScenGen to produce meaningful and scenario-aligned testing sequences. RQ4 examines the computational and temporal efficiency of test generation. Finally, RQ5 assesses the framework’s effectiveness in identifying real-world software bugs, demonstrating its practical utility. Together, these RQs provide a multi-dimensional understanding of the capability and robustness of ScenGen.
LLM Configuration. We employ the GPT-4V model (et al., 2024) as the LLM instance in ScenGen, leveraging its advanced visual understanding capabilities for reasoning about GUI layouts and scenario semantics. The model is accessed through the gpt-4-vision-preview API to enable direct visual processing of app screenshots. The parameter temperature is fixed at 0 to promote determinism and minimize stochastic response variation across runs. This setting ensures consistency in logical reasoning and decision outcomes while maintaining fairness and reproducibility in testing results.
@@
Scenario and App Selection. To establish realistic and representative testing targets, we conduct a systematic investigation of the Google Play Store to identify the most prevalent app categories based on download volume and the number of available apps. From these categories, we extract the most frequent and essential business functionalities as candidate testing scenarios. The test scenarios are created by three experienced software testing researchers, each with over five years of experience in mobile app testing and GUI automation. To reduce subjectivity, the researchers follow a structured scenario definition protocol: (1) each scenario is first described in natural language to represent a specific end-user goal; (2) all draft scenarios are reviewed by the research team to ensure coverage across different functional domains; and (3) overlapping or ambiguous scenarios are merged or refined through discussion. This collaborative design process ensures that the scenarios are representative of realistic user interactions while minimizing developer bias.
@@
As summarized in TABLE 1, ten representative testing scenarios are finalized, covering a range of functional complexities. The first column lists the scenario names, the second provides textual descriptions, and the third shows the number of apps associated with each. Based on these scenarios, we select a total of 92 apps (with 7 used in two scenarios) and construct 99 tasks (i.e., one specific scenario on one specific app). The full list of apps is available on our project website. To enhance coverage and realism, we include both open-source and commercial apps collected from multiple major platforms, including the Google Play Store and the Huawei AppGallery. Since our LLM performs best on English-language content, only apps with English interfaces are selected to eliminate confounding effects due to multilingual GUIs. The chosen apps span a wide range of functionalities and interaction complexities, ensuring the diversity and external validity of our evaluation.
@@
Baselines. We employ several baselines to facilitate fair and meaningful comparison across research questions. For RQ1 and RQ2, which focus on internal mechanisms of ScenGen, we use quantitative evaluations without external baselines. For RQ3 and RQ4, we compare against three state-of-the-art scenario-based or context-aware testing frameworks: GPTDroid (Liu et al., 2024a), ScenTest (Yu et al., 2024a), and ASOT [^1] Not an official name from the authors. (Su et al., 2025). These baselines represent recent advances in scenario-guided test generation. For RQ5, which examines bug detection effectiveness, we benchmark against four representative automated GUI testing tools: Monkey (mon, 2024) (random-based), Stoat (Su et al., 2017) (model-based), PIRLTest (Yu et al., 2024b) (learning-based), and ScenTest (Yu et al., 2024a) (scenario-based), covering a spectrum of exploration strategies. This selection ensures that performance is compared across both traditional and modern paradigms in automated GUI testing.
4.2. RQ1: Logical Decision-Making Effectiveness
RQ1 evaluates whether the Decider can accurately interpret the requirements of a testing scenario and generate appropriate next-step operations based on the current app GUI state. The goal is to assess both the logical soundness of initial decisions and the effectiveness of the self-correction mechanism when feedback from the environment necessitates revision. We use decision accuracy as the primary evaluation metric. To measure the impact of the self-correction mechanism, each decision and its potential subsequent correction are treated as a single decision unit. We then calculate the initial accuracy ($Acc_{1}$) and final accuracy ($Acc_{f}$) according to:
Table 2. Result of RQ1: Logical Decision-Making Effectiveness
Scenario N $n_{1}$ $n_{f}$ $Acc_{1}$ $Acc_{f}$ $\uparrow$
Email 39 38 39 97.44% 100.00% 2.63%
Music 31 31 31 100.00% 100.00% 0.00%
Note 71 64 66 90.14% 92.96% 3.13%
Photo 18 18 18 100.00% 100.00% 0.00%
Translation 51 48 48 94.12% 94.12% 0.00%
Shopping 38 36 36 94.74% 94.74% 0.00%
Login 42 40 41 95.24% 97.62% 2.50%
Calculation 57 57 57 100.00% 100.00% 0.00%
Weather 28 27 27 96.43% 96.43% 0.00%
Alarm-clock 33 33 33 100.00% 100.00% 0.00%
Sum/Avg 408 392 396 96.08% 97.06% 1.02%
Across the ten evaluated scenarios, the initial decision accuracy ranges from 90.14% to 100.00%, while the final accuracy ranges from 92.96% to 100%. On average, the initial accuracy ($Acc_{1}$) is 96.82%, and the final accuracy ($Acc_{f}$) increases to 97.59%, demonstrating that the self-correction mechanism yields a measurable improvement in decision reliability. In four scenarios, the Decider achieves perfect accuracy without correction, whereas in three other scenarios, accuracy improves noticeably after correction. This result confirms that the self-correction mechanism effectively compensates for transient decision errors caused by dynamic UI updates or ambiguous widget states.
4.3. RQ2: Widget Localization Effectiveness
Widget localization plays a crucial role in enabling accurate and reliable GUI interactions during automated testing. Unlike pure vision-based or text-based methods, ScenGen integrates the pixel-level precision of traditional computer vision (CV) algorithms with the semantic reasoning capabilities of multi-modal LLMs to handle visually complex and context-rich user interfaces. RQ2 aims to evaluate the effectiveness of this hybrid strategy in widget matching and localization, examining how the integration of CV techniques and LLM-driven understanding improves both accuracy and robustness in real-world testing environments.
Table 3. Result of RQ2: Widget Localization Effectiveness
Scenario N $n_{1}$ $n_{f}$ $Acc_{1}$ $Acc_{f}$ $\uparrow$
Email 33 28 32 84.85% 96.97% 14.29%
Music 22 12 22 54.55% 100.00% 83.33%
Note 53 44 52 83.02% 98.11% 18.18%
Photo 9 6 9 66.67% 100.00% 50.00%
Translation 35 34 34 97.14% 97.14% 0.00%
Shopping 31 26 30 83.87% 96.77% 15.38%
Login 35 32 33 91.43% 94.29% 3.13%
Calculation 44 34 44 77.27% 100.00% 29.41%
Weather 27 21 26 77.78% 96.30% 23.81%
Alarm-clock 23 21 23 91.30% 100.00% 9.52%
Sum/Avg 312 258 305 80.79% 97.96% 24.71%
The experimental results are summarized in TABLE 3. Across ten functional scenarios, the initial localization accuracy ($Acc_{1}$) ranges from 54.55% to 97.14%, while the final localization accuracy ($Acc_{f}$) after self-correction improves substantially, reaching between 94.29% and 100%. In four scenarios, the final accuracy achieves a perfect 100%. On average, the initial accuracy is 80.79%, and after correction, the average rises to 97.76%, with all scenarios exceeding 94% in the final stage.
Furthermore, qualitative analysis reveals that most initial localization errors occur in cases where widgets are visually occluded (e.g., drop-down or scrollable menus) or contain highly similar visual components (e.g., repeated icons). The self-correction mechanism successfully resolves these issues by leveraging LLM-based semantic reasoning to re-evaluate contextual cues and by refining coordinate-level predictions based on CV feedback. This cooperative process effectively mimics human testers’ behavior when re-targeting misclicked or ambiguous GUI elements. Overall, RQ2 demonstrates that combining computer vision algorithms with multi-modal LLM reasoning significantly enhances widget localization reliability in GUI testing. The self-correction mechanism further strengthens this process by dynamically adapting to interface changes and mitigating recognition errors. These results validate that ScenGen achieves precise, context-aware localization across diverse GUI layouts, ensuring a stable foundation for accurate and scenario-driven automated test execution.
4.4. RQ3: Scenario-based Test Generation Capability
@@
In RQ3, we evaluate the overall capability of ScenGen to automatically generate GUI tests that accurately and reliably complete predefined testing scenarios. This assessment aims to verify whether ScenGen can effectively understand, plan, and execute multi-step scenario-specific testing processes, thereby demonstrating its practical usability and reliability.
@@
Table 4. Result of RQ3: Scenario-based Test Generation Capability (in total of 99 tasks)
@@
The experimental results are summarized in TABLE 4. Across 99 total test generation tasks, ScenGen achieves the highest Coverage Rate (86.87%) and Success Rate (84.85%) among all compared methods. GPTDroid, ScenTest, and ASOT obtain Coverage Rates of 74.75%, 77.78%, and 84.85%, respectively, and Success Rates of 68.69%, 69.70%, and 73.74%. This comparison demonstrates that ScenGen surpasses all baselines in both comprehensive coverage and stable completion across diverse testing scenarios. Specifically, GPTDroid and ScenTest, though both scenario-oriented, rely primarily on text-based reasoning or pre-defined rules for test planning, leading to higher rates of premature or incomplete executions. ASOT achieves relatively strong results due to its adaptive scenario graph modeling, yet it lacks the semantic flexibility to handle unexpected GUI transitions or dynamically generated elements. In contrast, ScenGen benefits from its LLM-guided multi-agent design and context memory integration, allowing it to interpret GUI semantics, dynamically reason about progress within scenarios, and adapt decision-making to interface feedback.
@@
4.5. RQ4: Scenario-based Test Generation Efficiency
4.5.1. Token Consumption Analysis
Table 5. Result of RQ4: Token Overhead of Scenario-based Test Generation
Scenario $T_{1}$ $T_{2}$ $T_{3}$ $T_{4}$ $T_{5}$ $T_{6}$ $T_{total}$
Email 5586 8295 5281 13970 11823 833 45788
Music 3748 10197 4982 14729 8841 1963 44460
Note 6461 13639 6602 21667 12178 2573 63120
Photo 1486 3057 1791 6533 3571 746 17184
Translation 4645 6452 5640 11758 10296 359 39150
Shopping 8782 14461 12845 26102 19164 2757 84111
Login 8434 13263 7830 21329 18205 1253 70314
Calculation 5167 9184 6116 15603 12117 1940 50127
Weather 4522 11104 5474 12961 9419 1696 45176
Alarm-clock 3490 3794 3229 9120 8270 456 28359
Average 5232 9345 5979 15377 11388 1458 48779
4.5.2. Time Overhead Analysis
Table 6. Result of RQ4: Time Overhead of Scenario-based Test Generation (minute)
Scenario GPTDroid ScenTest ASOT ScenGen
Email 6.77 1.27 8.46 4.22
Music 7.13 1.24 7.63 3.37
Note 7.21 1.94 12.37 4.65
Photo 8.97 2.95 13.26 5.86
Translation 7.37 1.35 8.16 4.65
Shopping 6.58 2.16 11.69 4.81
Login 8.71 1.82 12.70 5.55
Calculation 7.36 2.10 10.68 3.15
Weather 6.78 1.30 11.54 5.50
Alarm-clock 7.93 1.05 7.35 5.39
Average 7.48 1.72 10.38 4.71
As shown in TABLE 6, the average time overhead of ScenGen (4.71 minutes) is higher than that of ScenTest (1.72 minutes) but significantly lower than that of ASOT (10.38 minutes) and GPTDroid (7.48 minutes). This demonstrates that ScenGen achieves a balanced trade-off between test generation efficiency and scenario reasoning capability. While its runtime exceeds ScenTest due to additional LLM interactions and vision-language inference, the increased computational effort translates directly into superior coverage and success rates, as demonstrated in RQ3.
RQ4 confirms that ScenGen achieves efficient and resource-aware scenario-based test generation. Its token consumption scales proportionally with scenario complexity, reflecting effective resource allocation across different reasoning stages. In terms of time overhead, ScenGen demonstrates moderate runtime with stable performance across diverse app scenarios, significantly faster than other advanced LLM-based frameworks while maintaining superior reliability and adaptability. These findings affirm that ScenGen provides a computationally efficient yet semantically powerful solution for real-world scenario-based GUI testing.
4.6. RQ5: Bug Detection Capability

### v3

@@ 2.1. Testing Scenario
In real-world app usage, end users interact with mobile apps through a series of GUI operations, such as tapping, swiping, or text input, to accomplish a specific goal. These operation sequences form logical patterns that can be abstracted into corresponding testing scenarios. For example, the “user login” scenario may include operations like entering credentials, tapping the login button, and verifying access to the home page. Understanding such scenarios is critical for testing app functionality from a user-centric perspective, as it enables testers to validate whether the app behaves as intended under realistic usage contexts. In the practice of manual GUI testing, developers and testers typically organize test cases around distinct testing scenarios. This organization allows them to design test cases more systematically, ensuring that all major functionalities and user flows are adequately exercised. Testing scenarios thus serve as a bridge between app functionality and user behavior, facilitating an intuitive and comprehensive verification of the app behavior.
@@ 3. Methodology
Specifically, the Observer perceives and interprets the app GUI by detecting widgets and extracting their semantic attributes, enabling the system to understand the current interface context. Building upon this perception, the Decider performs reasoning and planning, leveraging LLM capabilities to align GUI interactions with the target testing scenario and business logic. The Executor then executes the planned operations on the app, ensuring repeatability and traceability while decoupling action execution from decision-making to reduce the risk of cascading failures. Next, the Supervisor verifies whether the executed operations achieve the intended scenario objectives and provides feedback for correction when discrepancies occur. Finally, the Recorder records all operations, GUI states, and inspection outcomes into the shared context memory, preserving contextual data to support consistent and informed decision-making in subsequent iterations.
@@ 3.1. Context Memory
The context memory is designed to maintain continuity and coherence throughout the iterative testing process, ensuring that ScenGen can generate logically consistent scenario-based GUI tests. However, large language models (LLMs) are inherently stateless, lacking the ability to retain contextual information between interactions. This limitation poses a major challenge for multi-step GUI testing, where each decision depends on accumulated knowledge of previous actions, app states, and scenario progress. Without persistent context, the testing process risks producing inconsistent or redundant behaviors. To overcome this issue, ScenGen introduces a structured context memory inspired by human memory theory (Baddeley, 1992). The context memory acts as a central repository for storing and managing diverse types of information generated during testing, thereby supporting reasoning, decision-making, and feedback across agents. Specifically, it is divided into three components: long-term memory, which stores static information such as device configuration, app metadata, and testing scenarios; working memory, which maintains dynamic information including current testing goals, executed actions, and agent interactions; and short-term memory, which records transient GUI states and recognition results. This hierarchical organization enables seamless information flow across iterations, ensuring that testing decisions remain contextually grounded and progressively adaptive.
@@ 3.2. Observer
The Observer is responsible for perceiving the app’s GUI and constructing a structured representation of its current state to support subsequent decision-making and testing operations. However, accurately understanding GUI states presents a dual challenge: traditional computer vision (CV) techniques offer precise pixel-level detection but lack semantic comprehension, while large language models (LLMs) possess strong reasoning abilities yet struggle with fine-grained localization. To overcome this limitation, ScenGen adopts a hybrid perception design that combines traditional CV algorithms with multi-modal LLM reasoning. The CV component extracts and localizes widgets through visual analysis, while the LLM interprets their semantic roles and contextual relationships within the interface. This integration enables the Observer to achieve both spatial precision and semantic depth, producing a comprehensive, human-like understanding of the GUI state essential for accurate, scenario-driven test generation.
@@ 3.3. Decider
The Decider is responsible for determining the next operation during scenario-based GUI test generation, guiding the testing process toward fulfilling the target scenario. In ScenGen, test generation is modeled as an iterative decision-making process that begins from the app’s initial state and advances through a sequence of context-aware actions until the testing goal is achieved. However, this process presents two key challenges: first, accurately reasoning about the next action requires understanding both the current GUI state and the overall business logic of the target scenario; second, LLMs, while capable of semantic reasoning, may struggle to align their decisions with precise, executable GUI operations. To address these issues, ScenGen employs an LLM-guided decision mechanism that transforms GUI understanding and scenario semantics into concrete, stepwise operations. Each iteration evaluates the current app state, the history of executed actions, and the predefined scenario context to infer the most appropriate next operation through LLM reasoning. This design enables the Decider to bridge abstract scenario understanding and concrete GUI interactions, ensuring that every decision aligns logically with the scenario progression and contributes effectively to scenario completion. Every action decision is derived as an inference mapping generated by an LLM, formulated as follows:
@@
Compared to traditional LLMs, multi-modal LLMs can directly process GUI screenshots, enabling visual-semantic reasoning and action planning. Thus, ScenGen employs raw GUI images and utilizes multi-modal LLMs to make visual decisions directly. However, this design introduces a cognition–execution gap: while LLMs excel at reasoning about what should be done, they lack the ability to locate precise pixel-level widget positions required for action execution (Liu et al., 2024b). To address this, ScenGen integrates multi-modal reasoning with traditional CV-based widget localization. The decision process is thus divided into two stages, logical decision-making and widget localization, bridging high-level reasoning and low-level action feasibility.
@@ 3.3.2. Target Widget Location
After logical decision-making, the action intent and the target description are available, but precise pixel-level widget coordinates are still unknown. To enable actionable execution, the Decider performs widget localization, refining the decision into an executable operation:
@@ 3.5. Supervisor
The Supervisor is responsible for verifying the app state after each executed action, ensuring that the operation outcomes align with expectations before the next decision is made. Accurate verification is essential to maintain consistency between scenario intent and actual app behavior during iterative test generation. However, GUI state validation in mobile apps is inherently challenging due to dynamic page transitions, asynchronous loading behaviors, and diverse feedback mechanisms. Without robust verification, minor visual delays or hidden transitions may lead to misjudged states, erroneous decisions, or premature scenario termination. To address these challenges, the Supervisor employs a three-stage verification mechanism consisting of real-time loading verification, state transition verification, and test completion verification. This design allows the system to first ensure the app has reached a stable state, then confirm that state changes correspond to the expected outcomes, and finally determine whether the target testing scenario has been completed. By continuously validating and correcting deviations, the Supervisor provides a feedback loop that enhances system reliability, ensuring accurate and traceable progress throughout scenario-based test execution.
@@
Real-time Loading Verification. The execution of a test operation may temporarily place the app into an unstable state, such as a loading or transition phase triggered by a network request. In such cases, premature decision-making may lead to misinterpretation of the app state or the early triggering of self-correction. While a fixed waiting time could mitigate this issue, it is difficult to generalize and often harms testing efficiency. Instead, ScenGen leverages the visual understanding capability of a multi-modal LLM to dynamically determine whether the app has completed loading. By analyzing the current GUI image, the Supervisor detects loading indicators or transitional states and adjusts the waiting duration adaptively. This adaptive waiting strategy improves both the reliability and efficiency of the testing process. An example of this is shown in Fig. 4 (a).
@@ 4.1. Experiment Settings
  • RQ1: How effective is the scenario-based test generation of ScenGen?
  • RQ2: How effective is the bug detection capability of ScenGen?
  • RQ3: How effective is the widget localization of ScenGen?
  • RQ4: How effective is the logical decision-making of ScenGen?
  • RQ5: How efficient is the scenario-based test generation of ScenGen?
These research questions jointly provide a comprehensive assessment of ScenGen from multiple perspectives. Specifically, RQ1 investigates the overall quality of the generated tests, verifying the ability of ScenGen to produce meaningful and scenario-aligned testing sequences. RQ2 assesses the framework’s effectiveness in identifying real-world software bugs, demonstrating its practical utility. RQ3 measures the precision of widget localization conducted by the Observer, focusing on accurate visual grounding within GUI states. RQ4 evaluates the accuracy of logical decision-making by the Decider, assessing the LLM’s ability to produce semantically valid and contextually relevant actions. Finally, RQ5 examines the computational and temporal efficiency of test generation. Together, these RQs provide a multi-dimensional understanding of the capability and robustness of ScenGen.
LLM Configuration. We employ the GPT-4V model (et al., 2024) as the LLM instance in ScenGen, leveraging its advanced visual understanding capabilities for reasoning about GUI layouts and scenario semantics. The model is accessed through the gpt-4-vision-preview API, which enables direct visual processing of app screenshots. GPT-4V is chosen for its strong performance in multi-modal reasoning tasks, its ability to interpret complex visual contexts in combination with textual prompts, and its demonstrated robustness in structured decision-making. These characteristics make it particularly suitable for GUI testing, where accurate comprehension of widget semantics, layout hierarchy, and visual cues is essential for generating meaningful scenario-driven actions. The parameter temperature is fixed at 0 to promote determinism and minimize stochastic response variation across runs, ensuring consistency in logical reasoning and decision outcomes while maintaining fairness and reproducibility in testing results.
@@
Scenario and App Selection. To establish realistic and representative testing targets, we conduct an investigation of the Google Play Store to identify the most prevalent app categories based on download volume and the number of available apps. From these categories, we extract the most frequent and essential business functionalities as candidate testing scenarios. The test scenarios are created by three experienced software testing researchers, each with over five years of experience in mobile app testing and GUI automation. To reduce subjectivity, the researchers follow a structured scenario definition protocol: (1) each scenario is first described in natural language to represent a specific end-user goal; (2) all draft scenarios are reviewed by the research team to ensure coverage across different functional domains; and (3) overlapping or ambiguous scenarios are merged or refined through discussion. This collaborative design process ensures that the scenarios are representative of realistic user interactions while minimizing developer bias.
@@
As summarized in Table 1, ten representative testing scenarios are finalized, covering a range of functional complexities. The first column lists the scenario names, the second provides textual descriptions, and the third shows the number of apps associated with each. Based on these scenarios, we select a total of 92 apps (with 7 used in two scenarios) and construct 99 tasks (i.e., one specific scenario on one specific app). The full list of apps is available on our project website. To enhance coverage and realism, we include both open-source and commercial apps collected from multiple major platforms, including the Google Play Store and the Huawei AppGallery. Since our LLM performs best on English-language content, only apps with English interfaces are selected to eliminate confounding effects due to multilingual GUIs. The chosen apps span a wide range of functionalities and interaction complexities, ensuring the diversity and validity of our evaluation.
@@
Baselines. We employ several baselines to facilitate fair and meaningful comparison across research questions. For RQ1 and RQ5, we compare against three state-of-the-art scenario-based or context-aware testing frameworks: GPTDroid (Liu et al., 2024a), ScenTest (Yu et al., 2024a), and ASOT [^1] Not an official name from the authors. (Su et al., 2025). These baselines represent recent advances in scenario-guided test generation. For RQ2, which examines bug detection effectiveness, we benchmark against four representative automated GUI testing tools: Monkey (mon, 2024) (random-based), Stoat (Su et al., 2017) (model-based), PIRLTest (Yu et al., 2024b) (learning-based), and ScenTest (Yu et al., 2024a) (scenario-based), covering a spectrum of exploration strategies. This selection ensures that performance is compared across both traditional and modern paradigms in automated GUI testing. For RQ3 and RQ4, which focus on internal mechanisms of ScenGen, we use quantitative evaluations without external baselines.
4.2. RQ1: Scenario-based Test Generation Capability
@@
In RQ1, we evaluate the overall capability of ScenGen to automatically generate GUI tests that accurately and reliably complete predefined testing scenarios. This assessment aims to verify whether ScenGen can effectively understand, plan, and execute multi-step scenario-specific testing processes, thereby demonstrating its practical usability and reliability.
@@
Table 2. Result of RQ1: Scenario-based Test Generation Capability (in total of 99 tasks)
@@
The experimental results are summarized in Table 2. Across 99 total test generation tasks, ScenGen achieves the highest Coverage Rate (86.87%) and Success Rate (84.85%) among all compared methods. GPTDroid, ScenTest, and ASOT obtain Coverage Rates of 74.75%, 77.78%, and 84.85%, respectively, and Success Rates of 68.69%, 69.70%, and 73.74%. This comparison demonstrates that ScenGen surpasses all baselines in both comprehensive coverage and stable completion across diverse testing scenarios. Specifically, GPTDroid and ScenTest, though both scenario-oriented, rely primarily on text-based reasoning or pre-defined rules for test planning, leading to higher rates of premature or incomplete executions. ASOT achieves relatively strong results due to its adaptive scenario graph modeling, yet it lacks the semantic flexibility to handle unexpected GUI transitions or dynamically generated elements. In contrast, ScenGen benefits from its LLM-guided multi-agent design and context memory integration, allowing it to interpret GUI semantics and dynamically adapt decision-making to interface feedback.
@@
4.3. RQ2: Bug Detection Capability